# Macroeconomic Spillovers of Weather Shocks across U.S. States

### v2

@@ 5 Conclusions
We have investigated the effects of weather shocks on economic activity across U.S. states, with a particular focus on cross-border spillover effects. Using a GVAR framework, we have shown that severe weather shocks generally cause drops in state-level economic activity, with negative effects propagating to other regions through economic linkages between states. We estimate that an extreme weather event involving one climate region of the U.S. has the potential to reduce regional yearly growth rates by up to 1.3 percentage points, approximately. Because of indirect (second-round) effects, network relationships between states amplify the total impact that weather shocks have on the states where they occur.

### v3

@@ 6 Conclusions and policy implications
We have investigated the short-run effects of weather shocks on economic activity across U.S. states, with a particular focus on cross-border spillover effects. Using a GVAR framework, we have shown that severe weather shocks generally cause drops in state-level economic activity, with negative effects propagating to other regions through economic linkages between states. We estimate that an extreme weather event involving one climate region of the U.S. has the potential to reduce regional yearly growth rates by up to 1.8 percentage points, approximately. Because of indirect (second-round) effects, network relationships between states amplify the total impact that weather shocks have on the states where they occur.
Our results provide several implications from a policy perspective. Moreover, such implications concern both federal and local governments. On the one hand, the short-run negative effects on the economy of the region directly affected by an adverse weather shock suggest that local policymakers should plan in advance about the possible consequences of sudden increases in spending needs and, at the same time, declines in the revenues deriving from taxation. On the other hand, our results about economic spillovers reveal that all the other states (not just those geographically close to the area affected by the natural disaster) are not immune from serious adverse economic consequences. This evidence suggests the need for coordination of local policies to shield regional economies from severe weather shocks, but also for a direct and proactive intervention of the federal government to supervise and coordinate local policies, providing strategic actions for dealing with extreme natural events which may become more and more frequent in the future.